# Well-clipped cones behave themselves under all finite quotients, the cone conjecture under most

### v1

@@ 3.B. Non-examples and questions on well-clipped cones
the cone $\mathscr{C}$ is not locally rational polyhedral. In particular, this provides a subspace of $\mathfrak{S}$ containing $v_{k}$ of dimension $119-\dim\mathscr{A}_{k}+$ 1, thus $\dim\mathscr{A}_{k}\geq 10$.
@@
An important fact to keep in mind, mentioned by Totaro in [45, Section 3], is how Assumption (ii) of Definition 3.1 tends to fail in presence of canonical surface singularities. We present an example inspired by [45, Section 3].
@@ 3.C. Fundamental domains and well-clipped cones
By Lemma 3.13, we can apply [46, Theorem 1, Proposition 6] to the cone $\Pi$. We obtain that $w=1$, which concludes the proof of the disjoint interior property.
@@
    The nef cone of any smooth surface underlying a klt Calabi–Yau pair is perfectly clipped by Proposition 3.18 below and by the cone conjecture in dimension $2$ proven by Totaro [45].
@@
    The movable cone of a smooth projective hyperkähler manifold is perfectly clipped by the work of Markman [28, Theorem 6.18 (4+5), Lemma 6.23], notably thanks to the Hodge theoretic Torelli theorem [28, Theorem 1.3]. Please note the words of caution in [28, Paragraph following Definition 1.1, Caution 6.19] and the counterexamples [8, 35, 27], that convinced us to work on descending the movable (rather than the nef) cone conjecture and to take arithmetic subgroups in Definition 3.15 rather than the whole group $\mathrm{Aut}(\mathscr{A},V_{\mathbb{Z}})$ (compare with [36]).
@@
    The movable cone of a projective primitive symplectic variety $X$ with terminal $\mathbb{Q}$-factorial singularities is perfectly clipped. It is clear for $b_{2}(X)\leq 4$ by Item (1) above, and follows for $b_{2}(X)\geq 5$ from the work of Bakker-Lehn [2, Theorem 8.2] and Lehn-Mongardi-Pacienza [23, Theorem 5.12], using notably a Torelli theorem [23, Theorem 4.9].
@@
    The nef cone of $\mathbb{P}^{2}$ blown up at $9$ general points and the movable cone of $\mathbb{P}^{3}$ blown up at 8 very general points are both perfectly clipped, by Proposition 3.18 below and by [33] and [42] respectively.
@@
Our motivation to introduce these notions of neatly and perfectly clipped cones is the following proposition, which is inspired by the work of Ash [1, Chapter II], Sterk [43], Markman [28], Looijenga [25, Example 4.8], and Lehn–Mongardi–Pacienza [23]. One can view it as a characterization of those well-clipped cones which admit rational polyhedral fundamental domains under some group action.
@@
We start with the self-dual homogeneous cone $\mathscr{A}$. Let $\Gamma$ be an arithmetic subgroup of $\mathrm{Aut}(\mathscr{A})$ that contains all the orthogonal reflections $(\sigma_{i})_{i\in I}$ with respect to the hyperplanes $(H_{i})_{i\in I}$ cutting out $\mathscr{C}$. By the work of Vinberg on reduction theory and Siegel sets, and by [1, Chapter II, Theorem 4.1] and [25, Proposition 4.2], there is a rational polyhedral fundamental domain for the action of $\Gamma$ on the cone $\mathscr{A}^{+}$. By [25, Application 4.14], we can perform the following construction à la Dirichlet–Voronoi: If we choose a point $a\in\mathscr{A}\cap V_{\mathbb{Q}}$ with trivial stabilizer in $\Gamma$, the cone
@@
is a rational polyhedral fundamental domain for the action of $\Gamma$ on $\mathscr{A}^{+}$. From here on, we fix $a$ to be a point in $\mathscr{C}^{\circ}\cap V_{\mathbb{Q}}$ that has trivial stabilizer in $\Gamma$, and whose $\Gamma$-orbit avoids the hyperplanes $(H_{i})_{i\in I}$ used to cut out $\mathscr{C}$. (Its existence essentially follows from [25, Theorem 3.8], see [12, Page 8, Proof of Proposition 2.3] and [24, Page 13, Proof of Lemma 3.5, Last Paragraph] for more details.)
@@
for some subgroup $\Gamma_{\mathscr{C}}<\mathrm{Aut}(\mathscr{C}^{\circ},\mathscr{A},V_{\mathbb{Z}})$, and for a group $W$ such that non-trivial $W$-translates of $\mathscr{C}^{\circ}$ avoid $\mathscr{C}^{\circ}$. Note that for $\gamma\in\Gamma$, we have $\gamma(\mathscr{C}^{\circ})\cap\mathscr{C}^{\circ}\neq\varnothing$ if and only if $\gamma\in\Gamma_{\mathscr{C}}.$ So $\mathfrak{S}=\Gamma_{\mathscr{C}}$ is a subgroup of $\mathrm{Aut}(\mathscr{A},V_{\mathbb{Z}})$, and $\Pi$ transparently is a rational polyhedral fundamental domain for its action. Since $\Gamma_{\mathscr{C}}$ is contained in $\mathrm{Aut}(\mathscr{C}^{\circ},\mathscr{A},V_{\mathbb{Z}})$, there also is a rational polyhedral fundamental domain for the action of the larger group $\mathrm{Aut}(\mathscr{C}^{\circ},\mathscr{A},V_{\mathbb{Z}})$ on $\mathscr{C}^{+}$, see for instance [25, Proposition 4.1, Application 4.14]. The fact that $\Gamma_{\mathscr{C}}$ has finite index in $\mathrm{Aut}(\mathscr{C}^{\circ},\mathscr{A},V_{\mathbb{Z}})$ is then a consequence of the general fact [25, Proposition 4.6]. ∎
@@
We now assume Item (iii) and prove Item (i). By the work of Vinberg on reduction theory and Siegel sets, and by [1, Chapter II, Theorem 4.1] and [25, Proposition 4.2], there is a rational polyhedral fundamental domain for the action of the arithmetic group $\mathrm{Aut}(\mathscr{A},V_{\mathbb{Z}})$ on the cone $\mathscr{A}^{+}$. Thus and by [25, Proposition 4.6], the cone $\mathscr{C}$ is perfectly clipped if and only if there also is a rational polyhedral fundamental domain for the action of the group
@@ 4.A. Descent of self-dual homogeneous cones
Let $\mathscr{A}$ be a self-dual homogeneous cone in $V$. Let $G$ be a finite subgroup of $\mathrm{Aut}(\mathscr{A})$. The invariant cone $\mathscr{A}^{G}\subset V^{G}$ is self-dual homogeneous. Moreover, if $\mathscr{A}$ is compatible with a rational structure $V_{\mathbb{Q}}$ on $V$, then $\mathscr{A}^{G}$ is compatible with the rational structure $V_{\mathbb{Q}}^{G}$. Furthermore, if $\mathscr{A}$ is of hyperbolic type, then $\mathscr{A}^{G}$ is a halfline, the direct sum of two halflines, or a cone of hyperbolic type.
@@
Since $\mathscr{A}$ is stable by sum and $G$-invariant, the invariant cone $\mathscr{A}^{G}$ is clearly non-empty. By Lemma 2.6, we take an $\mathrm{Aut}(\mathscr{A},V_{\mathbb{Z}})$-invariant quadratic form $\mathrm{tr}$ on $V$ with respect to which $\mathscr{A}$ is self-dual. Since it is in particular $G$-invariant, its restriction to $V^{G}$ remains non-degenerate, and the cone $\mathscr{A}^{G}$ is clearly self-dual with respect to it.
@@
Note that $\mathscr{A}$ is compatible with the rational structure $V_{\mathbb{Q}}$ if and only if the corresponding Jordan operation $\circ$ (for a choice of $e\in\mathscr{A}\cap V_{\mathbb{Q}}$) restricts to an operation on $V_{\mathbb{Q}}$. This prves the “Moreover” part of this lemma.
Now, if $\mathscr{A}$ is of hyperbolic type, there is a hyperbolic quadratic form $q$ on $V$ such that $\partial\mathscr{A}\subset\{v\in V\mid q(v)=0\}$. The boundary of $\mathscr{A}^{G}$ is also contained in the zero locus of the restriction $q|_{V^{G}}$. If $\dim V^{G}\leq 2$, the cone $\mathscr{A}^{G}$ is a direct sum of one or two halflines, as wished. Assume now that $\dim V^{G}\geq 3$. If $q|_{V^{G}}$ is a hyperbolic quadratic form, then we are done. Otherwise, note that $q|_{V^{G}}$ cannot be negative definite: Thus, it must degenerate. Let $v\in V^{G}$ be such that $q(v,V^{G})=\{0\}$. Since $q$ is $G$-invariant, we have that
$|G|\cdot q(v,w)=q\left(v,\sum_{g\in G}g(w)\right)=0$
@@
The proof of Lemma 2.14 involves fixing a $G$-invariant element $e$ in $\mathscr{A}$. This is essentially the same choice as the one made by Monti–Quedo in [31, Page 13, Paragraph below Lemma 6.6] of a $G$-invariant polarization on the abelian variety with a $G$-action.
@@
Let $\mathscr{A}$ be a self-dual homogeneous cone in $V$. Let $G$ be a finite subgroup of $\mathrm{Aut}(\mathscr{A})$. Then, the centralizer $C_{\mathrm{Aut}(\mathscr{A})}(G)$ acts transitively on the cone $\mathscr{A}^{G}$.
@@
Let $e,a\in\mathscr{A}^{G}$. Consider the Jordan algebra $(V,e,\circ)$ associated to the cone $\mathscr{A}$ pointed at $e$ by Theorem 2.12. For $b\in V$, we denote by $L(b)$ the left-multiplication operator by $b$ with respect to the algebra law $\circ$. Following [10, Page 30, Before Proposition II.2.3], we say that an element $b\in V$ is invertible if there exists $y\in\mathbb{R}[b]$ such that $b\circ y=e$. By [10, Proposition II.3.1], if $b$ is invertible, then the quadratic representation $Q(b):=2L(b)^{2}-L(b^{2})$ is an invertible linear endomorphism of $V$. By [10, Proposition III.2.2], the operator $Q(b)$ also preserves the cone $\mathscr{A}$.
@@
see also [10, Page 48, Last line of Proof of Theorem III.2.1]. In particular, we can take $b\in V^{G}$ invertible element such that $a=b\circ b$. Then $Q(b)$ belongs to $\mathrm{Aut}(\mathscr{A})$ and sends $e$ to $a$. The fact that $Q(b)$ commutes with any $g\in G$ follows from the fact that both $b$ and $\circ$ are perserved by the $G$-action. ∎
@@ 4.B. Descent of well-clipped cones
Let $V$ be a finite dimensional $\mathbb{R}$-vector space, and let $G<\mathrm{GL}(V)$ be a finite subgroup. Then the invariant subspace $V^{G}$ is preserved by the action of the normalizer subgroup $N_{\mathrm{GL}(V)}(G)$. We denote by
@@
Let $\mathscr{A}$ be a self-dual homogeneous cone in $V=V_{\mathbb{Z}}\otimes\mathbb{R}$. Let $G$ be a finite subgroup of $\mathrm{Aut}(\mathscr{A},V_{\mathbb{Z}})$. Let $\mathscr{C}$ be a well-clipped cone in $\mathscr{A}$, cut out by hyperplanes $(H_{i})_{i\in I}$. Then there exist a subset $I^{*}\subset I$ and a self-dual homogeneous cone $\mathscr{B}\subset V^{G}=V^{G}_{\mathbb{Z}}\otimes\mathbb{R}$ with the same round part as $\mathscr{A}^{G}$, whose simplicial part is ruled by the simplical part of $\mathscr{A}^{G}$, such that
@@
The key of the proof is to extract orthogonal reflections on the invariant space $V^{G}$ out of a large Weyl group of orthogonal reflections on $V$, on which $G$ acts by permuting the (infinitely many) generators. Merely keeping the orthogonal reflections that are fixed by this action by permutation of $G$ is not enough: We need to build more. This difficulty is already mentioned in [36, 1.4], and the way we overcome it is quite similar to [36, 1.6]. However, we have to work in the framework of convex cones, which is not as nice as working in the Néron-Severi space of a K3 surface (the set-up of [36]) and bans us from using geometric tools like the Riemann–Roch and Noether formulas.
@@
where $({H_{i}}^{+})_{i\in I}$ is a minimal set of open halfspaces delimited by rational hyperplanes $H_{i}$ satisfying Assumptions (i), (ii), (iii). We choose the index set $I$ to be minimal is in Remark 3.2. For $i\in I$, we set $e_{i}\in H_{i,-}$ to be the generator of $H_{i}^{\perp}\cap V_{\mathbb{Z}}$, and we define $h_{i}:=H_{i}\cap V^{G}$, which is a hyperplane in $V^{G}$. Note that
$\varepsilon_{i}:=\sum_{g\in G}g(e_{i})$
@@
$q(\varepsilon_{i},\varepsilon_{k})=\sum_{g,g^{\prime}\in G}q(g(e_{i}),g^{\prime}(e_{k}))\geq 0,$
@@
(2) $\overset{\circ}{\mathscr{C}^{G}}=\mathscr{B}\cap\bigcap_{i\in j^{-1}(J_{1})}{h_{i}}^{+}\cap\bigcap_{\begin{subarray}{c}d\geq 3\\ i\in j^{-1}(J_{d})\end{subarray}}{h_{i}}^{+}.$
@@
(3) $\overset{\circ}{\mathscr{C}^{G}}=\mathscr{B}\cap\bigcap_{i\in I^{*}}{h_{i}}^{+},$
@@
Fix $i\in I^{*}$. Let $s_{i}:=-q(e_{i})\in\mathbb{Z}_{>0}$, let $G_{i}<G$ be the stabilizer of $e_{i}$, by $d:=|G|$ and $d_{i}:=|G_{i}|$. Let $\{g_{1},\ldots g_{d/d_{i}}\}$ be a set of representatives of the left-cosets $G/G_{i}$, with $g_{1}=1$. We take note that
$\eta_{i}:=\frac{1}{d_{i}}\varepsilon_{i}=e_{i}+\sum_{u=2}^{d/d_{i}}g_{u}(e_{i})$
still belongs to the invariant lattice $V_{\mathbb{Z}}^{G}$. Note that
(4) $-s_{i}+\sum_{u=2}^{d/d_{i}}q(e_{i},g_{u}(e_{i}))=q(e_{i},\eta_{i})=\frac{1}{dd_{i}}q(\varepsilon_{i})<0.$
@@
In particular, there is at most one $u\geq 2$ such that $q(e_{i},g_{u}(e_{i}))$ is non-zero, and then it equals $\frac{s_{i}}{2}$. Furthermore, $q(e_{i},\eta_{i})$ takes one of the two values $-s_{i},\frac{-s_{i}}{2}$.
Since $\eta_{i}$ spans the orthogonal of $h_{i}$, we can express the reflection
$\tau_{i}:x\in V^{G}\mapsto x-\frac{2q(e_{i},x)}{q(e_{i},\eta_{i})}\eta_{i}\in V^{G}.$
@@
This allows to decompose the orbit $G\cdot e_{i}$ into finitely many disjoint subsets $(B_{v})$ of size $1$ or $2$
@@
  • If all $B_{v}$ are of size $1$, we have $q(e_{i},\eta_{i})=q(e_{i})$, and
    $b|_{V^{G}}:x\in V^{G}\mapsto x-\frac{2q(e_{i},x)}{q(e_{i})}\eta_{i},$
    where we recall that $\eta_{i}$ is the sum of the orbit elements in $G\cdot e_{i}$ (taken once each), and use the fact that $q(e,x)=q(e_{i},x)$ for $x\in V^{G}$ and for any $e\in G\cdot e_{i}$.
@@
4.C. Normalizer, centralizer, and symmetries of the invariant cone

### v2

@@ 3.B. Non-examples and questions on well-clipped cones
the cone $\mathscr{C}$ is not locally rational polyhedral. In particular, this provides a subspace of $\mathfrak{S}$ containing $v_{k}$ of dimension $19-\dim\mathscr{A}_{k}+1$, thus $\dim\mathscr{A}_{k}\geq 10$.
@@
An important fact to keep in mind, mentioned by Totaro in [48, Section 3], is how Assumption (ii) of Definition 3.1 tends to fail in presence of canonical surface singularities. We present an example inspired by [48, Section 3].
@@ 3.C. Fundamental domains and well-clipped cones
By Lemma 3.13, we can apply [49, Theorem 1, Proposition 6] to the cone $\Pi$. We obtain that $w=1$, which concludes the proof of the disjoint interior property.
@@
    The nef cone of any smooth surface underlying a klt Calabi–Yau pair is perfectly clipped by Proposition 3.18 below and by the cone conjecture in dimension 22 proven by Totaro [48].
@@
    The movable cone of a smooth projective hyperkähler manifold is perfectly clipped by the work of Markman [30, Theorem 6.18 (4+5), Lemma 6.23], notably thanks to the Hodge theoretic Torelli theorem [30, Theorem 1.3]. Please note the words of caution in [30, Paragraph following Definition 1.1, Caution 6.19] and the counterexamples [9, 37, 29], that convinced us to work on descending the movable (rather than the nef) cone conjecture and to take arithmetic subgroups in Definition 3.15 rather than the whole group $\mathrm{Aut}(\mathscr{A},V_{\mathbb{Z}})$ (compare with [39]).
@@
    The movable cone of a projective primitive symplectic variety $X$ with terminal $\mathbb{Q}$-factorial singularities is perfectly clipped. It is clear for $b_{2}(X)\leq 4$ by Item (1) above, and follows for $b_{2}(X)\geq 5$ from the work of Bakker-Lehn [2, Theorem 8.2] and Lehn-Mongardi-Pacienza [25, Theorem 5.12], using notably a Torelli theorem [25, Theorem 4.9].
@@
    The nef cone of $\mathbb{P}^{2}$ blown up at 99 general points and the movable cone of $\mathbb{P}^{3}$ blown up at 8 very general points are both perfectly clipped, by Proposition 3.18 below and by [35] and [45] respectively.
@@
Our motivation to introduce these notions of neatly and perfectly clipped cones is the following proposition, which is inspired by the work of Ash [1, Chapter II], Sterk [46], Markman [30], Looijenga [27, Example 4.8], and Lehn–Mongardi–Pacienza [25]. One can view it as a characterization of those well-clipped cones which admit rational polyhedral fundamental domains under some group action.
@@
We start with the self-dual homogeneous cone $\mathscr{A}$. Let $\Gamma$ be an arithmetic subgroup of $\mathrm{Aut}(\mathscr{A})$ that contains all the orthogonal reflections $(\sigma_{i})_{i\in I}$ with respect to the hyperplanes $(H_{i})_{i\in I}$ cutting out $\mathscr{C}$. By the work of Vinberg on reduction theory and Siegel sets, and by [1, Chapter II, Theorem 4.1] and [27, Proposition 4.2], there is a rational polyhedral fundamental domain for the action of $\Gamma$ on the cone $\mathscr{A}^{+}$. By [27, Application 4.14], we can perform the following construction à la Dirichlet–Voronoi: If we choose a point $a\in\mathscr{A}\cap V_{\mathbb{Q}}$ with trivial stabilizer in $\Gamma$, the cone
@@
is a rational polyhedral fundamental domain for the action of $\Gamma$ on $\mathscr{A}^{+}$. From here on, we fix $a$ to be a point in $\mathscr{C}^{\circ}\cap V_{\mathbb{Q}}$ that has trivial stabilizer in $\Gamma$, and whose $\Gamma$-orbit avoids the hyperplanes $(H_{i})_{i\in I}$ used to cut out $\mathscr{C}$. (Its existence essentially follows from [27, Theorem 3.8], see [13, Page 8, Proof of Proposition 2.3] and [26, Page 13, Proof of Lemma 3.5, Last Paragraph] for more details.)
@@
for some subgroup $\Gamma_{\mathscr{C}}<\mathrm{Aut}(\mathscr{C}^{\circ},\mathscr{A},V_{\mathbb{Z}})$, and for a group $W$ such that non-trivial $W$-translates of $\mathscr{C}^{\circ}$ avoid $\mathscr{C}^{\circ}$. Note that for $\gamma\in\Gamma$, we have $\gamma(\mathscr{C}^{\circ})\cap\mathscr{C}^{\circ}\neq\varnothing$ if and only if $\gamma\in\Gamma_{\mathscr{C}}.$ So $\mathfrak{S}=\Gamma_{\mathscr{C}}$ is a subgroup of $\mathrm{Aut}(\mathscr{A},V_{\mathbb{Z}})$, and $\Pi$ transparently is a rational polyhedral fundamental domain for its action. Since $\Gamma_{\mathscr{C}}$ is contained in $\mathrm{Aut}(\mathscr{C}^{\circ},\mathscr{A},V_{\mathbb{Z}})$, there also is a rational polyhedral fundamental domain for the action of the larger group $\mathrm{Aut}(\mathscr{C}^{\circ},\mathscr{A},V_{\mathbb{Z}})$ on $\mathscr{C}^{+}$, see for instance [27, Proposition 4.1, Application 4.14]. The fact that $\Gamma_{\mathscr{C}}$ has finite index in $\mathrm{Aut}(\mathscr{C}^{\circ},\mathscr{A},V_{\mathbb{Z}})$ is then a consequence of the general fact [27, Proposition 4.6]. ∎
@@
We now assume Item (iii) and prove Item (i). By the work of Vinberg on reduction theory and Siegel sets, and by [1, Chapter II, Theorem 4.1] and [27, Proposition 4.2], there is a rational polyhedral fundamental domain for the action of the arithmetic group $\mathrm{Aut}(\mathscr{A},V_{\mathbb{Z}})$ on the cone $\mathscr{A}^{+}$. Thus and by [27, Proposition 4.6], the cone $\mathscr{C}$ is perfectly clipped if and only if there also is a rational polyhedral fundamental domain for the action of the group
@@ 4.A. Descent of self-dual homogeneous cones
Let $\mathscr{A}$ be a self-dual homogeneous cone in $V$. Let $G$ be a subgroup of $\mathrm{Aut}(\mathscr{A})$ acting with finite orbits. The invariant cone $\mathscr{A}^{G}\subset V^{G}$ is self-dual homogeneous. Moreover, if $\mathscr{A}$ is compatible with a rational structure $V_{\mathbb{Q}}$ on $V$, then $\mathscr{A}^{G}$ is compatible with the rational structure $V_{\mathbb{Q}}^{G}$. Furthermore, if $\mathscr{A}$ is of hyperbolic type, then $\mathscr{A}^{G}$ is a halfline, the direct sum of two halflines, or a cone of hyperbolic type.
@@
Since $\mathscr{A}$ is stable by sum and $G$-invariant, its contains the finite sum of all elements of a given $G$-orbit. Thus, the invariant cone $\mathscr{A}^{G}$ is clearly non-empty. By Lemma 2.6, we take an $\mathrm{Aut}(\mathscr{A},V_{\mathbb{Z}})$-invariant quadratic form $\mathrm{tr}$ on $V$ with respect to which $\mathscr{A}$ is self-dual. Since it is $G$-invariant, its restriction to $V^{G}$ remains positive definite and the cone $\mathscr{A}^{G}$ remains self-dual with respect to it.
@@
Note that $\mathscr{A}$ is compatible with the rational structure $V_{\mathbb{Q}}$ if and only if the corresponding Jordan operation $\circ$ (for a choice of $e\in\mathscr{A}\cap V_{\mathbb{Q}}$) restricts to an operation on $V_{\mathbb{Q}}$. This proves the “Moreover” part of this lemma.
Now, if $\mathscr{A}$ is of hyperbolic type, there is a hyperbolic quadratic form $q$ on $V$ such that $\partial\mathscr{A}\subset\{v\in V\mid q(v)=0\}$. The boundary of $\mathscr{A}^{G}$ is also contained in the zero locus of the restriction $q|_{V^{G}}$. If $\dim V^{G}\leq 2$, the cone $\mathscr{A}^{G}$ is a direct sum of one or two halflines, as wished. Assume now that $\dim V^{G}\geq 3$. If $q|_{V^{G}}$ is a hyperbolic quadratic form, then we are done. It is not negative definite since $q(e)>0$, thus it must degenerate. Let $v\in V^{G}$ be such that $q(v,V^{G})=\{0\}$. Since $q$ is $G$-invariant, we have that
$|G\cdot w|\cdot q(v,w)=q\left(v,\sum_{g\in G/{\rm Stab}(w)}g(w)\right)=0$
@@
The proof of Lemma 2.14 involves fixing a $G$-invariant element $e$ in $\mathscr{A}$. This is essentially the same choice as the one made by Monti–Quedo in [33, Page 13, Paragraph below Lemma 6.6] of a $G$-invariant polarization on the abelian variety with a $G$-action.
@@
Let $\mathscr{A}$ be a self-dual homogeneous cone in $V$. Let $G$ be a subgroup of $\mathrm{Aut}(\mathscr{A})$ acting with finite orbits. Then, the centralizer $C_{\mathrm{Aut}(\mathscr{A})}(G)$ acts transitively on the cone $\mathscr{A}^{G}$.
@@
Let $e,a\in\mathscr{A}^{G}$. Consider the Jordan algebra $(V,e,\circ)$ associated to the cone $\mathscr{A}$ pointed at $e$ by Theorem 2.12. For $b\in V$, we denote by $L(b)$ the left-multiplication operator by $b$ with respect to the algebra law $\circ$. Following [11, Page 30, Before Proposition II.2.3], we say that an element $b\in V$ is invertible if there exists $y\in\mathbb{R}[b]$ such that $b\circ y=e$. By [11, Proposition II.3.1], if $b$ is invertible, then the quadratic representation $Q(b):=2L(b)^{2}-L(b^{2})$ is an invertible linear endomorphism of $V$. By [11, Proposition III.2.2], the operator $Q(b)$ also preserves the cone $\mathscr{A}$.
@@
see also [11, Page 48, Last line of Proof of Theorem III.2.1]. In particular, we can take $b\in V^{G}$ invertible element such that $a=b\circ b$. Then $Q(b)$ belongs to $\mathrm{Aut}(\mathscr{A})$ and sends $e$ to $a$. The fact that $Q(b)$ commutes with any $g\in G$ follows from the fact that both $b$ and $\circ$ are perserved by the $G$-action. ∎
@@ 4.B. Descent of well-clipped cones
Let $V$ be a finite dimensional $\mathbb{R}$-vector space, and let $G<\mathrm{GL}(V)$ be a subgroup. Then the invariant subspace $V^{G}$ is preserved by the action of the normalizer subgroup $N_{\mathrm{GL}(V)}(G)$. We denote by
@@
Let $\mathscr{A}$ be a self-dual homogeneous cone in $V=V_{\mathbb{Z}}\otimes\mathbb{R}$. Let $G$ be a subgroup of $\mathrm{Aut}(\mathscr{A},V_{\mathbb{Z}})$ acting with finite orbits. Let $\mathscr{C}$ be a well-clipped cone in $\mathscr{A}$, cut out by hyperplanes $(H_{i})_{i\in I}$. Then there exist a subset $I^{*}\subset I$ and a self-dual homogeneous cone $\mathscr{B}\subset V^{G}=V^{G}_{\mathbb{Z}}\otimes\mathbb{R}$ with the same round part as $\mathscr{A}^{G}$, whose simplicial part is ruled by the simplical part of $\mathscr{A}^{G}$, such that
@@
The key of the proof is to extract orthogonal reflections on the invariant space $V^{G}$ out of a large Weyl group of orthogonal reflections on $V$, on which $G$ acts by permuting the (infinitely many) generators. Merely keeping the orthogonal reflections that are fixed by this action by permutation of $G$ is not enough: We need to build more. This difficulty is already mentioned in [39, 1.4], and the way we overcome it is quite similar to [39, 1.6]. However, we have to work in the framework of convex cones, which is not as nice as working in the Néron-Severi space of a K3 surface (the set-up of [39]) and bans us from using geometric tools like the Riemann–Roch and Noether formulas.
@@
where $({H_{i}}^{+})_{i\in I}$ is a minimal set of open halfspaces delimited by rational hyperplanes $H_{i}$ satisfying Assumptions (i), (ii), (iii). We choose the index set $I$ to be minimal is in Remark 3.2. For $i\in I$, we set $e_{i}\in H_{i,-}$ to be the generator of $H_{i}^{\perp}\cap V_{\mathbb{Z}}$, and we define $h_{i}:=H_{i}\cap V^{G}$, which is a hyperplane in $V^{G}$. Note that the finite sum
$\varepsilon_{i}:=\sum_{g\in G/{\rm Stab}(e_{i})}g(e_{i})$
@@
$q(\varepsilon_{i},\varepsilon_{k})=\sum_{\begin{subarray}{c}g\in G/{\rm Stab}(e_{i})\\ g^{\prime}\in G/{\rm Stab}(e_{k})\end{subarray}}q(g(e_{i}),g^{\prime}(e_{k}))\geq 0,$
@@
(2) $\overset{\circ}{\mathscr{C}^{G}}=\mathscr{B}\cap\bigcap_{i\in j^{-1}(J_{1})}{h_{i,+}}\cap\bigcap_{\begin{subarray}{c}d\geq 3\\ i\in j^{-1}(J_{d})\end{subarray}}h_{i,+}.$
@@
(3) $\overset{\circ}{\mathscr{C}^{G}}=\mathscr{B}\cap\bigcap_{i\in I^{*}}h_{i,+},$
@@
Fix $i\in I^{*}$. Let $s_{i}:=-q(e_{i})\in\mathbb{Z}_{>0}$. Let $\{g_{1},\ldots g_{d/d_{i}}\}$ be a set of representatives of the left-cosets $G/{\rm Stab}(e_{i})$, with $g_{1}=1$. We take note that
(4) $q(\varepsilon_{i})=|G\cdot e_{i}|\,q(e_{i},\varepsilon_{i})=|G\cdot e_{i}|\left(-s_{i}+\sum_{u=2}^{d/d_{i}}q(e_{i},g_{u}(e_{i}))\right).$
@@
In particular, there is at most one $u\geq 2$ such that $q(e_{i},g_{u}(e_{i}))$ is non-zero, and then it equals $\frac{s_{i}}{2}$. Furthermore, $q(e_{i},\varepsilon_{i})$ takes one of the two values $s_{i},\frac{-s_{i}}{2}$.
Since $\varepsilon_{i}$ spans the orthogonal of $h_{i}$, we can express the reflection
$\tau_{i}:x\in V^{G}\mapsto x-\frac{2q(e_{i},x)}{q(e_{i},\varepsilon_{i})}\varepsilon_{i}\in V^{G}.$
@@
This allows to decompose the orbit $G\cdot e_{i}$ into finitely many disjoint subsets $(B_{v})$ of size $1$ or 22
@@
  • If all $B_{v}$ are of size $1$, we have $q(e_{i},\varepsilon_{i})=q(e_{i})$, and
    $b|_{V^{G}}:x\in V^{G}\mapsto x-\frac{2q(e_{i},x)}{q(e_{i})}\varepsilon_{i},$
    where we recall that $\varepsilon_{i}$ is the sum of the orbit elements in $G\cdot e_{i}$ (taken once each), and use the fact that $q(e,x)=q(e_{i},x)$ for $x\in V^{G}$ and for any $e\in G\cdot e_{i}$.
@@
4.C. Centralizer and symmetries of the invariant cone